Gatti Manufacturing uses normal costing for its​ job-costing system, which has two​ direct-cost categories​ (direct materials and direct manufacturing​ labor) and one​ indirect-cost category​ (manufacturing overhead). The following information is obtained for 2017​:

Total manufacturing​ costs, $ 8,100,000$ times • Manufacturing overhead​ allocated, $3,600,000​(allocated at a rate of 250​% of direct manufacturing labor​ costs) times • Work-in-process inventory on January​ 2017​,$400,000 times • Cost of finished goods​ manufactured, $8,040,000

1.

Use information in the first two bullet points to calculate​ (a) direct manufacturing labor costs in

2017 and​ (b) cost of direct materials used in 2017.

2.

Calculate the ending​ work-in-process inventory on December​ 31,2017.
Answer #1

1) Direct manufacturing cost = Total manufacturing cost-Indirect manufacturing cost = 8100000-3600000 = 4500000

Direct labor = 3600000/2.5 = 1440000

Direct material = 4500000-1440000 = 3060000

2) Ending work in process = 400000+8100000-8040000 = $460000